OpenStudy (anonymous):

are these correct or incorrect? will give best response if you answer correctly

OpenStudy (anonymous):

1. is (z - 2)(z2+ 2z + 4)

OpenStudy (anonymous):

2.is correct

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what about 3 and 4?

OpenStudy (anonymous):

3. x^2 + 2x+ xy + 2y

OpenStudy (anonymous):

4. (x-7) (x+7) (x-2)

OpenStudy (anonymous):

oh sorry wait

OpenStudy (anonymous):

yeah i'm still confused on #3

OpenStudy (anonymous):

(x+y)(x+2) for 3

OpenStudy (anonymous):

sorry :p

OpenStudy (anonymous):

no need to be sorry haha, were these problems actually easy for you? like do you fully understand how to do them?

OpenStudy (anonymous):

1. c 2. c 3. d 4. d

OpenStudy (anonymous):

yeah do u want me to explain them to you?

OpenStudy (anonymous):

they're pretty easy

OpenStudy (anonymous):

okay yes please! so i can understand how to do similar problems in the future

OpenStudy (anonymous):

ok ill start with number three

OpenStudy (anonymous):

the right answer was (x+y) (x+2) I knew that by doing this: F:First x*x = x^2 O:Outer x*2= 2x I: Inner y*x = yx or xy it doesn't matter L: Last y * 2 = 2y put it all together x^2 + 2x+ xy + 2y

OpenStudy (anonymous):

oh okay I see how you got that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the rest are easy

OpenStudy (anonymous):

why dont you post them

OpenStudy (anonymous):

I posted them. Can you not see them?

OpenStudy (anonymous):

I think there was a lag because your responses weren't coming up but now they are.

OpenStudy (anonymous):

I'm not good at factoring sorry and I meant post them in a new question?

OpenStudy (anonymous):

Oh yeah, I guess that's a good idea. Thank you for your help with the previous four questions! Have a good rest of your day or night. :) I gave you best response.

OpenStudy (anonymous):

thnx and bye! tell me if you get all of them right!